Match the protons from the molecules below with the correct chemical shift on the 1H NMR spectra and the correct wavelengths on the FT-IR spectra.

102 100 92 84 82 78 76 74 72 70 68 64 60 1000 3500 2500 1500 Wavenumbers (cm-1)
NH O OH 1500 4000 3500 2500 2000 1000
8 106group133 1500 1000 4000 2500 Wavenumbers (ст.1)
102 100 92 84 82 78 76 74 72 70 68 64 60 1000 3500 2500 1500 Wavenumbers (cm-1)
group-117.102-epoxychalcone 1 1 Cridataichea3 23SP1 9 General DATE- 2019/04/09 TİHE 13:54 INSTRUM-spet P1 (18) I65536 P 400.13 폐_p. 8012.21 8 8 ppm)
Group_112 133 1 1 Crldata VChem323SP19 experinent product General 8 ATE2019/0402 IE 153 INSTRUMspec PULPROG "叩 NH O F1 (1H) OH I65536 SP400.13 S P8012.2 8 ppm)
NH O OH 1500 4000 3500 2500 2000 1000
8 106group133 1500 1000 4000 2500 Wavenumbers (ст.1)
Groun_ 1 12-133 1 1 С.VdataChem323SP 19 Erodures General ATE 2019/04 02 IM1531 ULPROG 2930 Fl (1H) I-65536 SP400.13 P8012.21 [ppm]
Answer #1

In the case of phenyl(3-phenyloxiran-2-yl)methanone:

FT-IR: epoxy:1280-1230 cm-1, carbonyl:1850, SP2 C-H strech:3050-3150, SP2 C-H bend: 650-1000, overtone patterns between 1660-2000, C=C: 1600-1480.

H1-NMR: 8.1-8.0 (2 H, multiplate), 7.67 (1H, triple triplate or Multiplate), 7.58 (2H, multiplate), 7.41 (5H, Multiplate), 4.2 (1H, doublate), 4.16 (1H, doublate)

2. 2-hydroxy-1,3-diphenyl-3-(phenylamino)propan-1-one:

FT-IR (cm-1):O-H: 3600-3500, N-H (2o): 1 band between 3300-3500, C=O 1715-1810, SP2 C-H bend: 650-1000, overtone patterns between 1660-2000, C=C: 1600-1480.

H1-NMR:7.97 (2H, M), 7.53 (2H, M), 7.63 (1H, M), 7.40 (2H, M), 7.23 (5H, M), 6.73 (3H, Chirality), 5.34 (1H, M), 4.52 (1H, chirality)